### v2

# DecoupledESC: Enhancing Emotional Support Generation via Strategy-Response Decoupled Preference Optimization

Abstract
Recent advances in Emotional Support Conversation (ESC) have improved emotional support generation by fine-tuning Large Language Models (LLMs) via Supervised Fine-Tuning (SFT). However, common psychological errors still persist. While Direct Preference Optimization (DPO) shows promise in reducing such errors through pairwise preference learning, its effectiveness in ESC tasks is limited by two key challenges: (1) Entangled data structure: Existing ESC data inherently entangles psychological strategies and response content, making it difficult to construct high-quality preference pairs; and (2) Optimization ambiguity: Applying vanilla DPO to such entangled pairwise data leads to ambiguous training objectives. To address these issues, we introduce Inferential Preference Mining (IPM) to construct high-quality preference data, forming the IPM-PrefDial dataset. Building upon this data, we propose a Decoupled ESC framework inspired by Gross’s Extended Process Model of Emotion Regulation, which decomposes the ESC task into two sequential subtasks: strategy planning and empathic response generation. Each was trained via SFT and subsequently enhanced by DPO to align with the psychological preference. Extensive experiments demonstrate that our Decoupled ESC framework outperforms baselines, reducing preference bias and improving response quality [^1] Our data and code are available at https://github.com/Zc0812/DecoupledESC..
Figure 1: Comparison from Vanilla-SFT to Vanilla-DPO to Decoupled-DPO. Vanilla-SFT lacks negative preference data, leading to high preference bias; Vanilla-DPO uses coupled preference data, causing potential negative optimization (regards PsNr, NsPr as pure negative samples); Decoupled-DPO decouples strategy and response, effectively reducing bias and psychological errors.
1 Introduction
Mental health is essential to well-being Prince et al. (2007), yet rising stress and fast-paced life have increased related issues Bor et al. (2014); Brundtland (2000); Paisley and McMahon (2001); Ma et al. (2023). According to WHO, 1/8 people suffer from mental disorders Organization (2022). Amid a shortage of professionals, this underscores the need for scalable solutions, where Large Language Models (LLMs) offer promising potential.
To enhance the performance of LLMs in Emotional Support Conversation (ESC), prior works Zhang et al. (2024); Chen et al. (2023) have constructed several large-scale, high-quality dialogue datasets and applied Supervised Fine-Tuning (SFT) to improve model responses. Among them, Liu et al. Liu et al. (2021) built the ESConv dataset based on Hill’s Helping Skills Theory Hill (1999) and filtered out FailedESConv dataset. The ESConv dataset follows a three-phase structure (Exploration → Comfort → Action) and includes eight types of support strategies, each paired with corresponding responses, details are provided in Appendix A and C.1. This structured design significantly enhances a model’s ability to generate empathetic dialogue.
Observation Currently, SFT has become the mainstream approach in the ESC field. However, we observe that models still frequently exhibit common psychological errors Raskin and Rogers (2005); Stebnicki (2007) during inference, which align with those identified in the FailedESConv dataset (Obs 1). In addition, Zhao et al. Zhao et al. (2025) found that SFT’s focus on single gold strategy-response pairs limits adaptability to nuanced contexts, weakening empathetic support. To mitigate this, they use Monte Carlo Tree Search (MCTS) to collect pairwise preference data linking strategies and responses, and apply Direct Preference Optimization (Vanilla-DPO) to guide the model in choosing appropriate strategies, thereby partially reducing preference bias and improving response quality.
Challenges However, as shown in Figure 1 and  4, our further analysis reveals that the limitations of current work lie not in the SFT or DPO training methods themselves, but rather in two overlooked challenges (Obs 2): (1) Entangled data structure: Existing ESC datasets heavily entangle psychological strategies with response content, making it difficult to construct high-quality preference pairs. For instance, penalizing responses with correct strategies but flawed content may degrade data quality. (2) Optimization ambiguity: Applying Vanilla-DPO directly to such entangled data can blur training objectives and even lead to negative optimization outcomes.
Approach To address these issues, we first introduce the Inferential Preference Mining (IPM) method, which automatically constructs preference samples decoupled from strategy-response. Specifically, we use dynamic data routing to route four types of psychological error samples identified from the SFT model’s inference data to the DPO training stage of either strategy planning or response generation, depending on the error type. These samples are then paired with human-annotated ground truth samples to form the Inferential Preference Mining Preference Dialogues (IPM-PrefDial) dataset, containing 2121k strategy preference pairs and 1111k response preference pairs. This dataset provides disentangled and high-quality supervision signals for two separate DPO models. Building on this, we propose a Decoupled ESC optimization framework (DecoupledESC), grounded in the Extended Process Model of Emotion Regulation (EPMER) Gross (2015), which divides emotion regulation into three sequential stages: (1) Identification → (2) Strategy Selection → (3) Implementation, details are provided in Appendix A.1. Accordingly, we explicitly split the ESC task into two subtasks: Strategy Planning (SP) and Response Generation (RG).
Results Across multiple evaluation metrics, our decoupled optimization framework significantly outperforms joint training baselines. It not only enhances the diversity of strategy selection but also improves response quality and empathy.
Contributions Our key contributions are summarized as follows:
  • We analyze common psychological errors in existing SFT paradigms and introduce Inferential Preference Mining (IPM) and Expert-Guided ICL Annotation to construct IPM-PrefDial, a decoupled strategy–response dataset.
  • We propose a Decoupled ESC framework, which explicitly splits the ESC task into two subtasks: Strategy Planning and Response Generation, effectively mitigating preference bias and enhancing response quality.
  • Extensive experiments show that our Decoupled ESC optimization framework significantly outperforms joint optimization baselines across multiple evaluation metrics.
Figure 2: Comparison of common psychological error type proportions among the FailedESConv dataset, Qwen-SFT inference results, and Llama-SFT inference results. Other Error refers to non-psychological errors.
Subset Qu RP RF Sd AR PS In Ot
GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$ GT Pred $\mathcal{B}$
1 110 135 1.41 40 25 0.57 43 38 0.82 39 50 1.43 85 50 0.55 77 91 1.29 26 19 0.69 80 92 1.25
2 104 122 1.32 35 18 0.51 41 43 1.11 48 42 0.90 74 56 0.78 73 95 1.46 28 20 0.76 97 104 1.15
3 97 120 1.42 29 30 1.04 47 37 0.79 37 40 1.13 79 54 0.66 83 95 1.24 35 20 0.57 93 104 1.16
4 93 135 1.78 27 13 0.45 40 32 0.81 40 42 1.19 89 49 0.51 77 110 1.68 32 18 0.57 102 101 1.02
Table 1: Analysis of Strategy Distribution and Preference Bias ($\mathcal{B}$) Across Four Randomly Sampled Subsets. This expanded layout clarifies the relationship between Ground Truth (GT), Predictions (Pred), and the resulting Bias for each strategy (defined in Appendix A.4).
2 Related Work
Emotional Support Conversation.
Emotional Support Conversation (ESC) aims to alleviate users’ emotional distress through empathetic and supportive responses. Liu et al. Liu et al. (2021) first introduced the concept and built the ESConv dataset with 8 support strategies, 1.3k dialogues. They also released the FailedESConv dataset, containing 196 failed dialogues. Subsequent studies improved ESC systems by enhancing data quality Sun et al. (2021); Qiu et al. (2024); Chen et al. (2023), adding external strategy planners Deng et al. (2024); He et al. (2024, 2025), and incorporating commonsense reasoning Tu et al. (2022); Deng et al. (2023); Luo and Yang (2024). Supervised Fine-Tuning (SFT) remains the dominant training paradigm with strong real-world performance (e.g., MeChat Qiu et al. (2024), SweetieChat Ye et al. (2025)). Recently, preference-based methods like Direct Preference Optimization (DPO) Rafailov et al. (2023) have emerged. Zhao et al. Zhao et al. (2025) introduced DPO with MCTS-based data to jointly optimize strategies and responses. However, the fixed coupling limited independent optimization and resulted in lower response quality.
Reinforcement Learning for LLM.
Reinforcement Learning (RL) was initially introduced into LLM training to align with human preferences Ouyang et al. (2022). This approach uses a reward model to guide the optimization of the policy model via the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017). Recently, the Group Relative Policy Optimization (GRPO) algorithm was proposed to enhance model reasoning capabilities Shao et al. (2024), which eliminates the need for a critic model by using within-group rewards as advantages. While these online reinforcement learning methods are effective, they suffer from high computational costs and reliance on accurate reward modeling. As a simpler offline optimization algorithm, DPO optimizes the policy model from pairwise preference data directly without the need for reward modeling. Due to its simplicity and effectiveness, DPO has achieved significant success across multiple domains, including mathematical reasoning, code generation, and recommendation systems Lai et al. (2024); Zhang et al. (2025a, b); Chen et al. (2024).
3 Preliminary Observations
To investigate the causes of low response quality in the ESC task, we analyzed outputs from six models: Base, SFT [^2] Trained on ESConv Liu et al. (2021) datasets.
github.com/thu-coai/Emotional-Support-Conversation
, and DPO versions of Qwen2.5-7B-Instruct Team (2024) and Llama3.1-8B-Instruct Dubey et al. (2024).
3.1 Preference Bias and Psy-Errors (Obs 1)
Current Base and SFT models (e.g., Qwen-Base, Qwen-SFT) show strong strategy preferences Kang et al. (2024), often overusing fixed strategies and failing to adapt to users’ emotional states. As shown in Figure 3, the Base and SFT models show different levels of divergence from the ground truth in strategy distributions.
Additionally, to verify that the observed strategy preference is an intrinsic model bias, not an artifact of sampling variance, we performed a validation on four random subsets of the test set. As shown in Table 1, the results reveal a consistent strategy bias across all subsets. Despite fluctuations in ground truth counts, the preference bias remained stable, strongly suggesting it is an intrinsic characteristic of the model’s decision-making process.
To further explore the impact of preference bias on response quality, we compared the outputs of Qwen-Base, Qwen-SFT, Llama-Base, and Llama-SFT with the FailedESConv dataset. As shown in Figure 2, common psychological errors Raskin and Rogers (2005); Stebnicki (2007) observed in the SFT-generated responses frequently aligned with those found in FailedESConv [^3] We have manually filtered out samples classified as FailedESConv due to non-psychological errors, such as incomplete or repetitive dialogues., including: (1) Strategy Mismatch, (2) Lack of Empathy, (3) Early Emotion Shift, (4) Template Response, (5) Emotion Misread. The definitions and corresponding examples are detailed in Appendix A.2.
Figure 3: Strategy Distribution across different models.
Although SFT reduces some errors, the empathy quality remains unsatisfactory. We argue that this stems from the SFT paradigm’s reliance on high-quality samples Zhao et al. (2025) without incorporating negative supervision signals from the FailedESConv dataset, failing to address bias in strategy selection and emotional understanding.
Figure 4: (a) Preference Bias and (b) Strategy Preference across Qwen and Llama models trained on different preference datasets.
3.2 Limitations of the DPO Method (Obs 2)
To address Obs 1, a natural approach is to treat filtered failures as negative signals and train with DPO. Prior work Zhao et al. (2025) adopted a vanilla DPO setup that jointly optimizes strategy-response pairs. However, as shown in Figure 1 and Figure 3, Vanilla-DPO relies heavily on the Question strategy and shows a strong preference for it, which fails to significantly reduce preference bias (see section Results).
To investigate the failure of Vanilla-DPO in aligning with human preferences, we conduct a controlled study. We split the preference data into two types: ① (PsPr, PsNr): where the preferred sample has both a positive strategy (Ps) and positive response (Pr), and the non-preferred sample has a positive strategy (Ps) but a negative response (Nr). ② (PsPr, NsNr): where the non-preferred sample contains both negative strategy (Ns) and negative response (Nr).
We train models using each dataset on Qwen and Llama, and evaluate them on preference bias and strategy preference. As shown in Figure 4 (a) and (b), models trained on consistently outperform those trained on . It reduces preference bias and better aligns with diverse strategies.
These results show that Vanilla-DPO training with entangled pairs like harms strategy learning. This reveals two issues in Vanilla-DPO:
  1. (1)
    Entangled data structure: The coupling between strategy and response complicates the construction of high-quality preference data, highlighting the need for more rigorous evaluation and filtering methods.
  2. (2)
    Optimization Ambiguity: Entangled strategy and response training lead to optimization ambiguity or even negative optimization: mislabeling PsNr as a negative sample leads to negative optimization on strategy learning, while NsPr harms response learning.
According to Hill’s Helping Skills Theory Hill (1999) and Gross’s Extended Process Model of Emotion Regulation (EPMER) Gross (2015), strategies should precede response generation and serve as its guidance. In essence, the two are decouplable. Inspired by this, we propose a decoupled modeling and staged optimization framework for ESC, which separates strategy planning from response generation, enabling more structured and targeted improvements in dialogue quality.
Criteria Total Assistant User
ESConv # Dialogues 1,040
# Utterances 29,526 14,763 14,763
Avg. Turns of Dialogue 28.40 14.20 14.20
Avg. Char of Utterance 95.85 112.17 79.54
Criteria Total Qwen Llama
IPM-PrefDial # Strategy Pref-Pairs 21,370 10,651 10,719
# Response Pref-Pairs 11,887 6,041 5,846
Avg. Char of Chosen 124.89 124.72 125.06
Avg. Char of Rejected 83.82 81.04 86.59
# Lack Emp. Response 4,371 2,288 2,083
# Emo. Shift Response 3,600 1,814 1,786
# Temp. Res. Response 3,916 1,939 1,977
Table 2: Statistics of the ESConv and IPM-PrefDial Datasets. Char: Character, Pref-Pairs: Preference Pairs.
4 Datasets
4.1 Preference Dataset Construction
Expert-Guided ICL Annotation.
Due to the large scale of the dataset, relying solely on human experts for psychological error annotation is highly time- and labor-intensive. Therefore, we adopt an In-Context Learning (ICL) Brown et al. (2020) approach to guide LLM-based classification. Specifically, we engaged 3 professional psychologists to annotate representative examples of 5 common psychological errors, along with detailed explanations. These expert-labeled instances were then used as ICL prompts, enabling the LLM to perform classification in alignment with expert standards.
Inferential Preference Mining.
The absence of failure-aware learning in standard SFT contributes to psychological errors observed in Obs 1. To address these issues, we propose Inferential Preference Mining (IPM) for collecting high-quality preference data, and an Expert-Guided ICL approach that leverages LLMs’ in-context learning to identify psychological errors.
Specifically, we first use the Qwen-SFT and Llama-SFT models to generate responses on the ESConv dataset. These responses are then paired with the corresponding gold responses to construct candidate preference pairs. To ensure consistency between the LLM and psychological expert annotators, we apply an Expert-Guided ICL Annotation process that prompts the LLM to filter out pairs exhibiting four common types of psychological errors. The remaining high-quality pairs constitute the IPM-PrefDial dataset.
More concretely, for the Strategy Planner (SP), we pair suboptimal strategies $s_{r}$ exhibiting the psychological error (1) Strategy Mismatch with gold strategies $s_{c}$ and context $c$ to form $D_{\text{SP-dpo}}$:
$D_{\text{SP-dpo}}=\left\{\left(c^{(i)},s_{c}^{(i)},s_{r}^{(i)}\right)\right\}_{i=1}^{|D_{\text{SP-dpo}}|}.$ (1)
For the Response Generator (RG), the SFT model produces multiple responses based on the gold strategy. Suboptimal responses $a_{r}$ exhibiting the psychological errors (2) Lack of Empathy, (3) Early Emotion Shift, or (4) Template Response are identified via Expert-Guided ICL and paired with gold responses $a_{c}$ to form $D_{\text{RG-dpo}}$:
$D_{\text{RG-dpo}}=\left\{\left(c^{(i)},s^{(i)},a_{c}^{(i)},a_{r}^{(i)}\right)\right\}_{i=1}^{|D_{\text{RG-dpo}}|}.$ (2)
Figure 5: Comparison between previous vanilla SFT training paradigm and our proposed Decoupled ESC framework. The Decoupled ESC first undergoes SFT initialization, followed by DPO training using the IPM-PrefDial dataset.
[…]
Qwen 2.5-7B- Instruct Chosen 3.82 3.52 3.20 2.95
Rejected 3.65 3.09 2.41 2.32
Improve ($\uparrow$) 4.66% 13.92% 32.78% 27.16%
Llama 3.1-8B- Instruct Chosen 3.99 3.74 3.33 3.09
Rejected 3.93 3.21 2.40 2.39
Improve ($\uparrow$) 1.53% 16.52% 38.75% 29.29%
Table 3: LLM-based evaluation scores for chosen and rejected responses across four dimensions.
4.2 Datasets Statistics
ESConv Dataset.
We employ ESConv dataset for SFT training ($D_{\text{sft}}$), which includes $1,040$ dialogues with an average of $14.2$ turns and $95.9$ characters per turn. Strategy distribution and temporal trends are shown in Table 8 and Figure 8 in the Appendix C.1.
IPM-PrefDial Dataset.
IPM-PrefDial dataset contains $21,370$ strategy preference pairs ($D_{\text{SP-dpo}}$) and $11,887$ response preference pairs ($D_{\text{RG-dpo}}$). In $D_{\text{RG-dpo}}$, chosen responses average $124.89$ characters, rejected ones $83.82$. Major rejection reasons include Lack of Empathy ($4,371$), Early Emotion Shift ($3,600$), and Template Response ($3,916$). Details are in Appendix C.2.
4.3 Datasets Quality
We evaluate the content quality of $100$ samples from $D_{\text{RG-dpo}}$ using gpt-4.1-mini-2025-04-14444https://openai.com/index/gpt-4-1. As shown in Table 3, chosen responses outperform rejected ones across four LLM-based metrics, with over a $30\%$ gain in Empathy.
[…]
D-1 B-1 F1 R-L Flu. Pro. Emp. Hel. $\mathcal{B}\downarrow$ $\mathcal{Q_{W}}\uparrow$ $\mathcal{Q}\uparrow$
Qwen2.5- 7B-Instruct Vanilla Base 93.50 9.75 14.92 12.59 3.55 2.53 1.89 1.38 2.17 8.41 8.06
+Direct-Refine 95.79 10.91 16.26 14.35 4.17 2.97 2.20 1.77 1.54 13.52 10.46
+Self-Refine 97.04 10.28 15.85 13.85 3.68 2.64 1.94 1.34 1.45 10.92 9.63
+Emotion CoT 97.17 10.61 16.07 14.06 3.95 2.70 2.50 1.51 1.87 6.89 6.63
SFT 90.93 15.61 20.99 17.78 3.30 2.61 2.29 2.12 0.31 24.89 20.27
DPO 88.13 16.23 21.24 18.03 3.47 2.67 2.36 2.23 0.30 22.25 18.97
Decoupled Base 97.55 10.97 16.33 14.19 3.92 2.71 2.17 1.38 1.92 13.96 12.07
SFT 91.37 16.69 22.15 18.76 3.93 2.72 2.40 2.11 0.27 26.94 21.37
DPO 89.84 17.73 22.86 19.31 3.99 2.90 2.54 2.02 0.22 27.09 21.77
Llama3.1- 8B-Instruct Vanilla Base 95.09 12.38 16.85 14.01 4.35 3.21 2.36 1.76 1.03 15.74 14.09
+Direct-Refine 90.07 11.36 14.97 12.79 3.35 2.82 2.16 1.35 1.72 12.12 9.98
+Self-Refine 87.18 10.72 14.26 12.20 3.53 2.95 2.40 1.45 1.68 13.93 12.00
+Emotion CoT 77.32 10.06 13.32 11.33 3.24 2.88 2.56 1.63 1.86 13.31 11.35
SFT 91.29 15.75 21.38 18.11 3.31 2.52 2.22 2.06 0.26 24.54 19.97
DPO 91.25 15.15 20.49 17.25 3.41 2.79 2.41 2.28 0.28 24.00 19.89
Decoupled Base 94.65 12.67 16.70 14.01 4.24 3.24 2.34 1.66 1.62 7.54 7.67
SFT 91.51 16.97 22.42 19.12 3.87 2.74 2.39 1.95 0.23 26.03 21.36
DPO 90.35 17.50 22.59 19.16 3.81 2.73 2.64 2.17 0.15 27.10 22.94
Table 4: Comparison of models under different optimization paradigms and training methods. The best score is in-bold, while the second best score is underlined. $\uparrow$ means a higher score is better whereas $\downarrow$ is exactly the opposite.
5 Methodology
5.1 Decoupled ESC Framework
To address the issue raised in Obs 1 and Obs 2, that vanilla training of strategy planning and response generation can lead to negative optimization, hindering the reduction of preference bias and the improvement of response quality. As shown in Figure 5, we propose a Decoupled ESC optimization framework, inspired by the Extended Process Model of Emotion Regulation (EPMER) Gross (2015), which divides emotion regulation into three sequential stages: Identification, Strategy Selection, and Implementation. We decouple the ESC generation process into two independent subtasks: Strategy Planning and Response Generation. This enables more stable and controllable training for each.
Specifically, we adopt a decoupled two-stage modeling framework: a Strategy Planner selects a optimal strategy based on the dialog history $c_{t}=(u_{0},a_{0},\ldots,u_{t-1},a_{t-1},u_{t})$, where $u$ and $a$ denote user and assistant utterances, respectively. The strategy is generated as $s_{t}\sim\text{LLM}_{\text{SP}}(s\mid c_{t})$. Then, a Response Generator generates an empathic reply conditioned on both the selected strategy and the dialog context: $a_{t}\sim\text{LLM}_{\text{RG}}(a\mid c_{t},s_{t})$.
5.2 Decoupled-SFT and Decoupled-DPO
Decoupled-SFT.
To optimize the performance of the Strategy Planner and Response Generator, we first initialize these two modules using the SFT method to endow them with the capabilities for strategy planning and empathic response generation. Specifically, based on real dialogues from the ESC dataset, we constructed a turn-level training dataset $D_{\text{sft}}=\left\{\left(c^{(i)},s^{(i)},a^{(i)}\right)\right\}_{i=1}^{|D_{\text{sft}}|}$. The two modules are then fine-tuned separately using SFT: (1) Strategy Planner: Using the dialogue context $c$ and the supporter’s response strategy $s$, we perform turn-level training to minimize the loss function:
$\mathcal{L}_{\text{SP-sft}}=-\mathbb{E}_{(c,s)\sim D_{\text{sft}}}\left[\log\mathrm{LLM}_{\text{SP}}(s|c)\right].$ (3)
(2) Response Generator: Given the context $c$, strategy $s$, and response $a$, minimizing the loss:
$\mathcal{L}_{\text{RG-sft}}=-\mathbb{E}_{(c,s,a)\sim D_{\text{sft}}}\left[\log\mathrm{LLM}_{\text{RG}}(a|c,s)\right].$ (4)
Decoupled-DPO.
To reduce psychological errors, we further optimize the Strategy Planner and Response Generator using the offline RL method (DPO). Based on the preference dataset IPM-PrefDial, which includes $D_{\text{SP-dpo}}$ and $D_{\text{RG-dpo}}$, we separately train both modules to enhance strategy selection and response generation. For the Strategy Planner, we apply DPO on $D_{\text{SP-dpo}}$ to encourage preference for gold strategies and reduce bias toward suboptimal ones. The loss function is defined as:
$\mathcal{L}_{\text{SP-dpo}}=-\mathbb{E}_{(c,s_{c},s_{r})\sim D_{\text{SP-dpo}}}\left[\log\sigma\Bigg(\beta\log\frac{\pi_{\theta}(s_{c}|c)}{\pi_{\text{ref}}(s_{c}|c)}\right$.
$\left.-\beta\log\frac{\pi_{\theta}(s_{r}|c)}{\pi_{\text{ref}}(s_{r}|c)}\Bigg)\right],$ (5)
For the Response Generator, we apply DPO on $D_{\text{RG-dpo}}$ to improve response quality and empathy, with the loss function defined as:
[…]
$\left.\left.-\beta\log\frac{\pi_{\theta}(a_{r}|c,s)}{\pi_{\text{ref}}(a_{r}|c,s)}\right)\right].$ (6)
where $\pi_{\theta}$ denotes the model being optimized, and $\pi_{\text{ref}}$ denotes the reference model after SFT.
6 Experiments
In this section, we conduct extensive experiments to address the following research questions:
  • RQ1: What are the performance differences between DPO and SFT in bias mitigation and response generation within a coupled framework?
  • RQ2: What specific advantages does decoupling strategy planning and response generation bring to ESC tasks?
  • RQ3: In the Decoupled ESC framework, how do SFT and DPO respectively affect the model’s bias and generation quality?
  • RQ4: To what extent can the Decoupled ESC framework effectively reduce psychological errors?
6.1 Experimental Setup
Backbones.
We conducted experiments using two LLMs: Qwen2.5-7B-Instruct Team (2024) and Llama3.1-8B-Instruct Dubey et al. (2024).
Baselines.
We compare vanilla coupled models (Base, SFT, DPO) with prompt-optimization baselines such as Direct-Refine, Self-Refine Madaan et al. (2023), and Emotional CoT Wei et al. (2022).
Datasets.
The ESConv Liu et al. (2021) dataset is split into train, valid, and test sets in an 8:1:1 ratio, with the training set used for SFT. The IPM-PrefDial dataset is used for DPO training.
Evaluation Metrics.
We evaluate model performance using the following metrics: (1) Automatic Metrics, including BLEU-1 (B-1) Papineni et al. (2002), Distinct-1 (D-1) Li et al. (2015), F1-score (F1), and ROUGE-L (R-L) Lin (2004); (2) LLM-based Metrics, including Fluency (Flu.), Professionalism (Pro.), Empathy (Emp.), and Helpfulness (Hel.). All metrics are rated on a 5-point Likert scale Joshi et al. (2015); (3) Strategy Metrics, including preference bias ($\mathcal{B}$Kang et al. (2024) and strategy prediction accuracy (weighted-F1 $Qw$ and Macro-F1 $Q$). Detailed definitions of the evaluation metrics, the prompt, and the Bias calculation formula are provided in Appendix E.
Implementation Details.
For SFT, we train for 3 epochs using a batch size of 32 and a learning rate of 1e-5. DPO is trained for 1 epoch under the same settings. All experiments are conducted on 4×24GB RTX 4090 GPUs. For LLM-based evaluation, we randomly sample 100 test instances and evaluate them using gpt-4.1-mini-2025-04-1444footnotemark: 4 alongside 3 psychology experts. Additional implementation details are provided in Appendix D.1.
[…]
Vanilla-SFT 15.75 18.11 3.31 2.52 2.22 2.06 0.26
Decoupled-SFT 16.97 19.12 3.87 2.74 2.39 1.95 0.23
Vanilla-DPO 15.15 17.25 3.41 2.79 2.41 2.28 0.28
Decoupled-DPO 17.50 19.16 3.81 2.73 2.64 2.17 0.15
Table 5: Results of SFT and DPO under Vanilla and Decoupled Paradigms for Llama3.1-8B-Instruct.
6.2 Experimental Results
Vanilla-DPO vs. Vanilla-SFT (RQ1).
As shown in Table 4, under the vanilla setting, DPO consistently outperforms SFT on LLM-based metrics for both Qwen and Llama, indicating enhanced response quality.
However, Strategy-Metrics evaluation shows mixed results: DPO slightly reduces bias ($\mathcal{B}$) on Qwen but increases bias and lowers accuracy on Llama. We attribute this to DPO’s sensitivity to noisy data in the vanilla setting—conflicting optimization signals between strategy and content can cause negative transfer (see section Obs 2). We further compare the impact of different preference data on coupled models in Appendix B.
Decoupled vs. Vanilla (RQ2).
As shown in Tables 4 and 5, decoupled models outperform vanilla ones on most metrics. Notably, the preference bias of Decoupled-DPO models drops to $0.22$ (Qwen) and $0.15$ (Llama), much better than Vanilla-DPO and Vanilla-SFT. This shows decoupled optimization effectively reduces preference bias.
Human Evaluation
[…]
Figure 6: Comparison of Human-Evaluated Win Rates for Decoupled-DPO (Llama) and Vanilla-DPO (Llama).
We attribute this to the decoupled framework. Its strategy planning module, as shown by Kang et al. Kang et al. (2024), acts as an external planner that helps reduce strategy preference bias. It also avoids the optimization conflict in vanilla training and simplifies preference data construction.
Decoupled-DPO vs. Decoupled-SFT (RQ3).
As shown in Tables 4 and 5, DPO brings greater improvements over SFT in the decoupled setting, with bias reduced from $0.27$ to $0.22$ for Qwen, and from $0.23$ to $0.15$ for Llama. In contrast, the vanilla framework yields minimal bias reduction from SFT to DPO, highlighting the stronger synergy between decoupling and DPO training. Table 6 further shows that, given ground-truth strategies, the DPO-trained Response Generator consistently outperforms its SFT counterpart across all LLM-based metrics.
[…]
Qwen2.5- 7B-Instruct 3.66 3.02 2.51 2.37
3.77 3.26 2.75 2.67
Llama3.1- 8B-Instruct 3.67 3.18 2.71 2.52
3.94 3.44 2.90 2.73
Table 6: Ablation study on Response Generator. ✓ indicates that the method or data is used in training, while ✗ indicates it is not. GT: Ground Truth Strategy.
Figure 7: Proportion of psy-error types in the response of Qwen and Llama under the decoupled framework.
Decoupled-DPO on Psychological Errors (RQ4).
To verify the effectiveness of Decoupled-DPO in improving response quality, we adopt the same error categorization approach as Obs 1. By contrasting Figures 2 and 7, we observe that Decoupled-DPO achieves the highest proportion of No Error cases at 2727%, outperforming Qwen-SFT and Llama-SFT by an average of 77%. This demonstrates that Decoupled-DPO effectively reduces common psychological errors and improves overall response quality.
7 Conclusion
In this paper, we propose a Decoupled ESC framework that separates strategy planning from empathetic response generation, enabling targeted optimization and avoiding mutual interference. Extensive experiments demonstrate that our Decoupled ESC framework significantly outperforms joint optimization baselines, effectively reducing preference bias and improving response quality in Emotional Support Conversation tasks.
Limitations
Our study, while demonstrating promising results, has several limitations that suggest avenues for future research. The annotation of psychological errors relied on an in-context learning approach to manage the resource-intensive nature of expert labeling; while efficient, this method’s classifications may not be perfectly consistent with expert judgment. Furthermore, our experiments were constrained by computational resources to models up to 9B parameters, so the efficacy of our method on much larger models (e.g., 70B) requires future validation. The generalizability of our decoupled framework also warrants broader investigation, as it was primarily tested with DPO and should be assessed with other algorithms like KTO, SimPO, and IPO. Finally, this work is confined to textual analysis, and a key future direction is to extend our framework to incorporate multimodal inputs such as speech and facial expressions for more holistic emotional support.
Ethics Statement
Data Usage Agreement
This research utilizes the ESConv and FailedESConv dataset Liu et al. (2021), which has been obtained with proper authorization and in compliance with data usage agreements. We ensure that all data used in this study is handled responsibly and in accordance with ethical standards, respecting the privacy and confidentiality of individuals involved. All necessary agreements and permissions for the use of this dataset have been signed, ensuring full compliance with data protection regulations.
Model Usage Policy
It should be noted that while the model demonstrates certain capabilities in psychological support tasks, its strategies cannot encompass the full range of approaches and techniques used in real-life professional counseling. Given the diversity of users’ emotional states and circumstances, the model’s responses may not always align with professional standards and may unintentionally affect users’ emotional well-being. Therefore, this model is intended for academic research only and is not recommended for commercial use. Caution is advised when using it beyond research settings, and it should not be applied to real-world counseling without professional supervision.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62293554, U2336212), “Pioneer” and “Leading Goose” R&D Program of Zhejiang (2024C01073), Ningbo Innovation “Yongjiang 2035” Key Research and Development Programme (2024Z292), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
Appendix A Definitions
A.1 Definitions of Gross’s Extended Process Model of Emotion Regulation
The Extended Process Model of Emotion Regulation (EPMER), proposed by Gross in 2015 Gross (2015), refines earlier models by conceptualizing emotion regulation as a temporally ordered process comprising 3 core stages:
  1. 1.
    Identification: Individuals assess whether an emotional response needs to be regulated based on situational goals and personal relevance.
  2. 2.
    Selection: A regulation strategy is chosen from available options, guided by the expected outcome of regulating the emotion.
  3. 3.
    Implementation: The selected strategy is carried out and monitored.
A.2 Definitions of Psychological Errors
Under the guidance of 3 psychological experts and based on psychological literature Raskin and Rogers (2005); Stebnicki (2007), we identified common errors frequently made by psychological experts in real-world therapy sessions. These were categorized into 5 types of empathy-related psychological errors [^5] All definitions of psychological errors were reviewed by 3 psychological experts.:
  • Strategy Mismatch: Selecting a strategy inappropriate for the user’s emotional state or context.
  • Lack of Empathy: Failing to recognize or validate the user’s emotional experience.
  • Early Emotion Shift: Prematurely changing the emotional tone before acknowledging the user’s current state.
  • Template Response: Relying on generic or scripted expressions lacking personalization.
  • Emotion Misread: Misinterpreting the user’s emotional cues, leading to unaligned responses.
Figures 11, 12, 13, and 14 illustrate representative examples of the first four error types, drawn from rejected responses in the IPM-PrefDial dataset.
A.3 Definitions of Counseling Stages
Liu et al. Liu et al. (2021) developed a three-stage counseling framework based on Hill’s Helping Skills Theory Hill (1999).
  1. 1.
    Exploration: Explore to identify the help-seeker’s problem.
  2. 2.
    Comforting: Comfort the help-seeker by expressing empathy and understanding.
  3. 3.
    Action: Assist the help-seeker in solving their problems.
Although most cases in our dataset follow the counseling sequence of (1) Exploration → (2) Comforting → (3) Action, some cases are adjusted based on the help-seeker’s specific situation.
[…]
PsPr NsNr PsNr NsPr D-1 B-1 F1 R-L Flu. Pro. Emp. Hel. $\mathcal{B}\downarrow$ $\mathcal{Q_{W}}\uparrow$ $\mathcal{Q}\uparrow$
Qwen2.5- 7B-Instruct 89.38 15.93 20.94 17.75 3.31 2.64 2.34 2.15 0.26 25.60 21.69
89.45 15.73 20.80 17.50 3.50 2.73 2.53 2.25 0.30 22.81 18.76
90.83 15.32 20.78 17.53 3.19 2.54 2.17 2.13 0.29 22.91 18.63
88.13 16.23 21.24 18.03 3.47 2.67 2.36 2.23 0.30 22.25 18.97
Llama3.1- 8B-Instruct 90.72 16.08 21.41 18.07 3.45 2.69 2.38 2.22 0.22 26.69 21.76
91.19 15.92 21.30 17.92 3.48 2.68 2.45 2.26 0.29 23.82 19.71
91.19 15.92 21.45 18.01 3.49 2.65 2.22 2.15 0.22 25.20 21.07
91.25 15.15 20.49 17.25 3.41 2.79 2.41 2.28 0.28 24.00 19.89
Table 7: Comparison of coupled models trained with different preference data. ✓ denotes that the training set contains this type of data, while ✗ denotes its absence in the training set. The best score is in-bold, while the second best score is underlined.
A.4 Definitions of Strategies
The strategies and its definitions in this study align with Liu et al. Liu et al. (2021) and follow Hill’s Helping Skills Theory Hill (1999).
  • Question (Qu): Asking for information related to the problem to help the help-seeker articulate the issues that they face. Open-ended questions are best, and closed questions can be used to get specific information.
  • Restatement or Paraphrasing (RP): A simple, more concise rephrasing of the help-seeker’s statements that could help them see their situation more clearly.
  • Reflection of Feelings (RF): Articulate and describe the help-seeker’s feelings.
  • Self-disclosure (Sd): Divulge similar experiences that you have had or emotions that you share with the help-seeker to express your empathy.
  • Affirmation and Reassurance (AR): Affirm the help-seeker’s strengths, motivation, and capabilities and provide reassurance and encouragement.
  • Providing Suggestions (PS): Provide suggestions about how to change, but be careful to not overstep and tell them what to do.
  • Information (In): Provide useful information to the help-seeker, for example with data, facts, opinions, resources, or by answering questions.
  • Others (Ot): Exchange pleasantries and use other support strategies that do not fall into the above categories.
Appendix B Analysis of Coupled Model Training Results
To further analyze the effects of varying preference data on coupled models, we evaluated coupled models trained with different preference data across multiple metrics, as shown in Table 7. The results indicate that the model trained with the suboptimal-content dataset (row 2, 6) significantly outperforms the model trained with the suboptimal-strategy dataset (row 3, 7) in terms of LLM-based metrics, while the reverse holds for strategy metrics. Additionally, it is notable that both the Vanilla-DPO model (row 4, 8) and the model trained with (PsPr, NsNr) data (row 1, 5) fail to achieve optimal performance across the two metric types. This further demonstrates that the coupled model has two optimization objectives, and it is not possible to achieve optimal performance on both objectives by fully utilizing the preference data. This indicates the effectiveness of the decoupled ESC framework.
Figure 8: Strategy distribution across dialogue stages in ESConv Dataset.
Figure 9: Strategy distribution in IPM-PrefDial Dataset.
Figure 10: Psychological Errors Distribution in Rejected Responses: Qwen and Llama.
Appendix C Datasets Details
C.1 ESConv and FailedESConv Datasets
Table 8 presents the number and proportion of support strategies in the ESConv dataset, while Figure 8 illustrates the distribution of these strategies across different dialogue stages. Figure 16 illustrates the prompt we use to classify the psychological errors in the FailedESConv dataset as well as the response content of Qwen-SFT and Llama-SFT.
Categories Number Proportion
Support Strategies Question (Qu) 3,060 20.73%
Resta. or Parap. (RP) 857 5.81%
Reflection (RF) 1,146 7.76%
Self-disclosure (Sd) 1,387 9.40%
Affir. & Reass. (AR) 2,288 15.50%
Suggestions (PS) 2,373 16.07%
Information (In) 989 6.70%
Others (Ot) 2,663 18.04%
Overall 14,763 100.00%
Table 8: Distribution of support strategies used in ESConv Dataset.
C.2 IPM-PrefDial Dataset
Figure 9 compares the distribution of support strategies in the Chosen and Rejected samples within the preference datasets of Qwen and Llama. Figure 10 further presents the count and proportion of psychological errors found in the rejected responses of these datasets. In addition, Figures 11, 12, 13, and 14 illustrate examples from the IPM-PrefDial dataset, covering both strategy preference and response preference data. Each example includes the dialogue context, as well as the chosen and rejected responses.
C.3 Prompts for Data Filter
Figure 15 presents the prompt we use to filter and select high-quality preference datasets, which effectively filters and identifies data that meets the required standards.
Backbone Model lr beta
Qwen2.5- 7B-Instruct Vanilla-dpo 7e-7 0.2
SP-dpo 5e-8 0.5
RG-dpo 7e-7 0.2
Llama3.1- 8B-Instruct Vanilla-dpo 5e-7 0.2
SP-dpo 8e-8 0.5
RG-dpo 3e-7 0.2
Table 9: Detailed training hyperparameters used in dpo.
Appendix D Implementation Details
D.1 Experiment Details
We employ Qwen2.5-7B-Instruct Team (2024) and Llama3.1-8B-Instruct Dubey et al. (2024) as our base models. All training procedures are implemented using the Llama-Factory framework Zheng et al. (2024) with LoRA fine-tuning Hu et al. (2022), where the alpha and rank are set to 16, and the dropout rate is 0.05. For SFT training, we trained the models for 3 epochs with the learning rate of 1e-5 and the batch size of 32. For DPO training, the batch size is 32 and the epoch is set to 1. We use vLLM Kwon et al. (2023) to accelerate the inference. All experiments are conducted on 4 NVIDIA RTX 4090 GPUs. More detailed hyperparameter settings for DPO are presented in Table 9.
D.2 Baselines
Direct-Refine. A straightforward self-optimization approach where the model directly revises its initial response to improve quality, without relying on external input or intermediate reasoning.
Self-Refine. Following Madaan et al. Madaan et al. (2023), this method involves two stages: the model first generates self-feedback on its initial response, then refines the output based on that feedback, promoting internal reflection and correction.
Emotional CoT. Extending Chain-of-Thought (CoT) prompting Wei et al. (2022), this method first elicits the user’s emotional state through intermediate reasoning, which then guides strategy planning and response generation.
Appendix E Details of Evaluation
E.1 Strategy Metrics
According to Kang et al. (2024), the strategy preference is calculated by the following formula.
$p^{\prime}_{i}=\frac{\sum_{j}\left(w_{ij}p_{j}\right)/\left(p_{i}+p_{j}\right)}{\sum_{j}w_{ji}/\left(p_{i}+p_{j}\right)},$ (7)
where $w_{ij}$ denotes the frequency count of the model predicting strategy $i$ given that the ground-truth strategy is $j$. All of the strategy preferences $p_{i}$ are initialized as 1 and updated through iteration of the preference bias.
The strategy preference bias $\mathcal{B}$ is computed from the strategy preference $p_{i}$ as follows:
$\mathcal{B}=\sqrt{\frac{\sum_{i=1}^{N}(p_{i}-\bar{p})^{2}}{N}},$ (8)
where $\bar{p}$ denotes the average strategy preference.
E.2 LLM Metrics Criteria
Table 10 summarizes the LLM evaluation metrics, including Fluency, Professionalism, Empathy, and Helpfulness, along with their descriptions, evaluation criteria, and scoring scales. All metrics are rated on a 5-point Likert scale Joshi et al. (2015). Specifically, Fluency and Empathy are adapted from the ESC-Eval framework Zhao et al. (2024), Professionalism is guided by the CPsyCoun framework Zhang et al. (2024), and Helpfulness is derived from the SoulChat evaluation setup Chen et al. (2023).
E.3 Prompt for LLM Metrics
Figures 171819, and 20 present the prompts used for LLM-based evaluation of Fluency, Professionalism, Empathy, and Helpfulness, respectively. Each prompt explicitly defines the role of the LLM as a judge and outlines the corresponding evaluation criteria. To minimize potential bias, the prompts are carefully designed to avoid revealing model names or being influenced by text length.
E.4 Human Evaluation
To complement the LLM-based evaluation and enhance the credibility of our results, we conducted a human evaluation with 3 licensed psychology experts on 100 samples generated by our Decoupled-DPO model based on the Llama backbone. In addition, we performed a pairwise win/loss comparison between Decoupled-DPO and Vanilla-DPO (both using Llama) on another set of 100 samples, with inter-rater agreement again measured by Fleiss’ Kappa ($\kappa$) Fleiss and Cohen (1973). As shown in Figure 6, Decoupled-DPO consistently outperforms Vanilla-DPO across all four LLM-based metrics. The evaluator agreement, measured by $\kappa$, further supports the reliability of the human judgments.
Figure 11: Example of psychological errors under Strategy Mismatch in rejected response from IPM-PrefDial.
Figure 12: Example of psychological errors under Lack of Empathy in rejected response from IPM-PrefDial.
Figure 13: Example of psychological errors under Early Emotion Shift in rejected response from IPM-PrefDial.
Figure 14: Example of psychological errors under Template Response in rejected response from IPM-PrefDial.
Figure 15: Prompt for Filtering High-Quality Preference Data in IPM-PrefDial.
Figure 16: Prompt for Classifying Psychological Errors in FailedESConv dataset, Qwen-SFT, and Llama-SFT Outputs.
Dimension Description Criterion Score
Fluency Fluency evaluates whether language expression is natural, coherent, and comprehensible. 1.1 Incoherent or difficult to understand; contains grammar or logic issues. 0
1.2 Unclear expression; user may struggle to grasp the meaning. 1
1.3 Some parts are confusing, though the main point can be inferred. 2
1.4 Mostly clear and coherent with minor ambiguities. 3
1.5 Fluent and well-structured; logically organized and easy to follow. 4
1.6 Concise and impactful language; precise and elegant communication that conveys ideas efficiently. 5
Professionalism Professionalism evaluates whether the model demonstrates psychological knowledge, follows ethical principles, and avoids misleading or inappropriate advice. 2.1 Contains harmful, misleading, or clearly inappropriate content that may violate ethical or psychological guidelines. 0
2.2 Shows serious misunderstanding or misuse of psychological concepts, or provides inappropriate advice. 1
2.3 Minor factual inaccuracies or advice that lacks evidence, but does not pose direct harm. 2
2.4 No major errors; advice is acceptable and somewhat aligned with psychological principles. 3
2.5 Demonstrates solid understanding of psychological concepts and appropriate intervention techniques. 4
2.6 Highly professional, reflects strong psychological insight, maintains boundaries, and communicates in a grounded, ethical manner. 5
Empathy Empathy evaluates whether the model genuinely understands the user’s emotions, expresses care, and provides emotional support. 3.1 Contains statements that may harm the user emotionally or lead to a negative emotional trajectory. 0
3.2 Fails to provide emotional comfort or assist the user in analyzing their problems. 1
3.3 Either lacks emotional comfort or fails to support problem analysis. 2
3.4 No significant issues, but empathy and analysis remain surface-level. 3
3.5 Demonstrates a warm, human-like tone—like a friend—offering both emotional relief and analytical support. 4
3.6 Deep emotional insight with sincere and stable empathy, conveyed through attentive and flexible language. 5
Helpfulness Helpfulness evaluates the effectiveness of an AI assistant’s suggestions by considering both the number of recommendations provided per interaction and the relevance or usefulness of each suggestion in addressing the user’s question. 4.1 Irrelevant, misleading, or potentially harmful suggestions. 0
4.2 Ineffective or generic advice that does not respond to the user’s needs. 1
4.3 Weakly relevant suggestions with limited practical value. 2
4.4 Somewhat helpful; suggestions are relevant and usable. 3
4.5 Clear and practical advice that aligns well with the user’s issue. 4
4.6 Highly insightful, tailored, and actionable suggestions that offer strong guidance and value. 5
Table 10: LLM Evaluation Metrics and Corresponding Score Criterion.
Figure 17: Prompt for Evaluating Fluency of Generated Responses via LLMs.
Figure 18: Prompt for Evaluating Professionalism of Generated Responses via LLMs.
Figure 19: Prompt for Evaluating Empathy of Generated Responses via LLMs.
Figure 20: Prompt for Evaluating Helpfulness of Generated Responses via LLMs.